# Guided Query Refinement: Multimodal Hybrid Retrieval with Test-Time Optimization

### v1

@@ Abstract
Multimodal encoders have pushed the boundaries of visual document retrieval, matching textual query tokens directly to image patches and achieving state-of-the-art performance on public benchmarks. Recent models relying on this paradigm have massively scaled the sizes of their query and document representations, presenting obstacles to deployment and scalability in real-world pipelines. Furthermore, purely vision-centric approaches may be constrained by the inherent modality gap still exhibited by modern vision-language models. In this work, we connect these challenges to the paradigm of hybrid retrieval, investigating whether a lightweight dense text retriever can enhance a stronger vision-centric model. Existing hybrid methods, which rely on coarse-grained fusion of ranks or scores, fail to exploit the rich interactions within each model’s representation space. To address this, we introduce Guided Query Refinement (GQR), a novel test-time optimization method that refines a primary retriever’s query embedding using guidance from a complementary retriever’s scores. Through extensive experiments on visual document retrieval benchmarks, we demonstrate that GQR allows vision-centric models to match the performance of models with significantly larger representations, while being up to 14x faster and requiring 54x less memory. Our findings show that GQR effectively pushes the Pareto frontier for performance and efficiency in multimodal retrieval. We release our code at this GitHub repository.
@@ 1 Introduction
Visual document retrieval is the task of returning relevant documents – typically PDFs containing figures, tables, and other visual elements – in response to a textual query (Mathew et al., 2021b; a; Li et al., 2024; Zhu et al., 2022; Faysse et al., 2025). To tackle this task, neural retrieval pipelines often follow a text-centric approach, relying on OCR or vision-language models to convert source documents into textual chunks, and then constructing an index using semantic text encoders (Karpukhin et al., 2020; Tanaka et al., 2021). An alternative, vision-centric, approach relies instead on multimodal encoder models. Building on the ColBERT (Khattab & Zaharia, 2020) late-interaction approach, ColPali-based (Faysse et al., 2025) encoders operate directly on image patches, and yield multi-vector embedding representations of images and queries.
@@
While this approach achieves state-of-the-art results on public benchmarks of visual document retrieval [^1] https://huggingface.co/blog/manu/vidore-v2 (Macé et al., 2025), open challenges within this paradigm remain. First, to pursue state-of-the-art performance, recent late-interaction multimodal retrievers [^2] Henceforth, we use retrievers and encoders interchangeably. massively scale the length and dimensionality of query and document representations. This can incur a substantial latency and storage overhead, hindering the ability to provide an efficient and scalable solution. For example, Llama-NemoRetriever-Colembed-3B represents each document page with 10 MB of memory (Xu et al., 2025), three orders of magnitude more than single-vector dense retrievers (Table 3). Secondly, a vision-centric approach for matching textual queries to textually rich documents may be limited by the substantial modality gap (Clavié & Brand, 2025; Li et al., 2025; Role et al., 2025) exhibited by modern vision-language models. These gaps motivate exploring complementary approaches for improving the performance of ColPali-based encoders.
@@
An early concept in the application of neural retrievers has been that of hybrid retrieval (dos Santos et al., 2015; Kuzi et al., 2020), where the outputs of different retrievers are aggregated at the level of ranks or query-document similarity scores (see Figure 1) to obtain the final list of retrieved documents. Hybrid retrieval (or hybrid search) most commonly refers to the combination of a neural semantic text retriever with a sparse lexical representation (e.g., BM25). More broadly, it reflects the notion that models relying on different types of representations can capture complementary aspects of the data, and boost the overall system performance.
In this work, we seek to connect these two threads, testing whether the paradigm of hybrid retrieval can complement modern multimodal encoders. Specifically, we aim to leverage the low latency and small storage footprint of dense text retrievers, and the uni-modal signal they provide, along with the benefits of ColPali-like vision-centric architectures.
@@
Standard hybrid retrieval methods rely on a rather coarse-grained view of the perspective of each retriever – they cannot utilize the rich query-document interactions within the model representation space (Figure 1). Aiming to harness this untapped potential for hybrid retrieval, in this work we propose Guided Query Refinement (GQR), a novel approach for aggregating retriever outputs. Given a query at test time, GQR iteratively optimizes the query representations of a primary retriever with gradient descent, using similarity scores from a complementary retriever. The refined query representation softly incorporates the complementary retriever’s signal, remaining subject to the query–document interactions in the primary retriever space. This updated query embedding is then used to score the documents and return an updated document list. Notably, GQR is architecture-agnostic and can be applied across single- and multi-vector retrievers.
@@ 2 Methods
We begin this section by outlining prominent methods at each level (§2.1, §2.2). We explain that while early representation-level aggregation could be desired due to its richness, it is difficult to achieve in practice. We then present our method, Guided Query Refinement (GQR), which lies between the levels of scores and of representations (§2.3).
@@ 2.1 Ranking-Level Aggregation
RRF (Cormack et al., 2009) combines ranked lists by weighting each item based on the reciprocal of its rank. The RRF constant $\kappa>0$ dampens the impact of very high ranks and controls how much credit is given to mid-list occurrences.
@@ 2.2 Score-Level Aggregation
Score-level aggregation operates one step deeper than ranking aggregation, operating on the real-valued similarity scores $s_{m}(q,d_{i})$ between the query and documents. To ensure that the scores of different retrievers are in the same scale and range, the common practice (Bruch et al., 2023) is to first apply a normalization function $N_{m}$ – yielding $\tilde{s}_{m}(q,d_{i})$ – and then aggregate across retrievers:
@@
More generally, the above methods can be viewed as a weighted aggregation, where the examples above are the uniform case, with each retriever assigned a weight $\alpha=1/M$ (for two retrievers, $\alpha=0.5$ for each). Given a development set, these weights can be fit (Bruch et al., 2023). Here, for $M=2$, the two retrievers are assigned relative weights $\alpha$ and $1-\alpha$, yielding the parameterized variants Average Ranking - Tuned, RRF - Tuned, Score Aggregation (Min-Max) - Tuned, and Score Aggregation (SoftMax) - Tuned.
@@ 2.3 Guided Query Refinement
Representation-level information carries the richest potential for effective aggregation. Embedding-level projections that align representational spaces are used extensively in modern vision–language systems to combine visual and textual inputs (Radford et al., 2021; Jia et al., 2021; Li et al., 2021; 2022). At test time, however, operating directly on representations is hindered by heterogeneity: encoders may use a single vector or many vectors per document and query, and they operate within differing dimensionalities and scales. Thus, with strict latency and memory budgets and without access to supervision, aggregation at the this level is not trivial. [^4] Concatenating single-vector embeddings is feasible, yet under dot-product scoring this reduces to an unnormalized sum of separate scores, and does not enable interaction between the spaces.
Our goal in this work is to exploit the rich information in query and document representations while remaining architecture-agnostic, lightweight, and practical. To this end, we propose Guided Query Refinement (GQR, Algorithm 1), a novel method for combining the outputs of two retrievers – a primary retriever $m_{1}$ and a complementary retriever $m_{2}$. GQR refines $m_{1}$’s query representation based on the signal of $m_{2}$’s scores. Our approach is inspired by query optimization methods that rely on pseudo-relevance feedback from a stronger cross-encoder at test time (Yu et al., 2021; Sung et al., 2023; Gangi Reddy et al., 2025). Here, instead of relying on a heavy cross-encoder, we utilize feedback from a lightweight bi-encoder, whose performance can be on par or even weaker than the primary encoder.
@@
We denote the initial query embedding of $m_{1}$ by $z^{(0)}=e_{1}^{q}$, and we update it at each step $t$, $z^{(t)}$ for $T$ steps. At step $t$, the consensus distribution of $m_{1}$ and $m_{2}$ is defined as
@@
Here, $\mathrm{KL}$ is the Kullback–Leibler divergence (Equation 9). Minimizing it pushes $p_{1}(\cdot\mid z^{(t)})$ to place higher probability where the consensus $p_{\mathrm{avg}}^{(t)}$ does, and to reduce probability where the consensus is low, aligning $m_{1}$ with the joint signal.
We apply a gradient step on the query representation with step size $\alpha$ [^5] We define GQR with gradient descent for simplicity, but in practice we found Adam (Kingma & Ba, 2015) to perform better and use it as the optimizer.,
@@
Gradients are computed through $m_{1}$’s scoring function, so the update remains tied to the interactions between the query and the documents in that representation space. We then compute the final scores from retriever $m_{1}$,
@@
Algorithm 1 Guided Query Refinement (GQR)
[…]
6:for $t=0$ to T1T-1 do
11:end for
[…]
14:return $\mathcal{R}(q)$
@@ 3.1 Setup
Visual document retrieval (Mathew et al., 2021b; a; Li et al., 2024; Zhu et al., 2022; Faysse et al., 2025) assumes a corpus of documents, that contain visual elements such as charts, images, and tables, and a set of document-grounded textual queries. The goal is to retrieve the most relevant documents for each query. We conduct experiments on ViDoRe 1 (Faysse et al., 2025) and ViDoRe 2 (Macé et al., 2025), which are established benchmarks for this task. Corpus documents are embedded by encoder models, either directly from page images, or following ingestion of document pages into text (see Appendix D).
@@
We evaluate a diverse pool of multimodal and textual state-of-the-art retrieval models. The Colpali-based set includes three encoders: Colnomic-Embed-Multimodal-7B (Team, 2025b), Jina-Embeddings-v4 (Günther et al., 2025), and Llama-NemoRetriever-Colembed-3B (Xu et al., 2025). The set of text models includes Linq-Embed-Mistral (Choi et al., 2024) and Qwen3-Embedding-4B (Zhang et al., 2025), as well as Jina-Embeddings-v4 in its multi-vector textual configuration. This yields $3$ text-based models and $3$ image-based models in total. See Table 3 for details on the models.
@@
We use $\mathrm{NDCG}@5$ as the primary metric for our evaluations. We also report $\mathrm{Recall}@5$ in Appendix E. For each ColPali-based vision-centric model, we test each of the text-centric models as the complementary retriever used for GQR. This yields $9$ GQR configurations in total, $3$ per ColPali-based model. We also evaluate 8 different hybrid baseline methods for each vision-text model pair. We tune GQR and the weighted hybrid methods on a development set for each dataset (see Appendix D for implementation details).
3.2 Results - Performance
@@
Comparison to model baselines.
@@
Table 1 reports GQR against the corresponding models on ViDoRe 2. The first block lists the text-only models, which average between $46.8$ for Qwen and $55.3$ for Linq. In each subsequent block, a ColPali-based retriever is fixed, and we show its score alongside the GQR variants, with deltas computed relative to the primary retriever. For Colnomic-7B, the average score rises from $60.3$ to $63.1$ with query refinement from Jina (text) $(+2.8)$ and to $62.8$ with Linq-Embed $(+2.5)$. For Llama-Nemo, the strongest model on ViDoRe 2 to date, GQR improves the average from $63.0$ to $65.2$ with Linq-Embed $(+2.2)$ and to $64.2$ with Jina (text) $(+1.2)$. Notably, the text models clearly underperform the ColPali-based retrievers, yet with GQR the complementary signal they provide boosts performance. This is clearest in the Llama-Nemo versus Qwen3 setting, where GQR delivers a small gain of $\uparrow$+0.3+0.3, despite a $16.2$ point gap in base $\mathrm{NDCG}@5$ (Llama-Nemo $63.0$ vs. Qwen $46.8$). Subsets where GQR harms performance are rare for Colnomic-7B and Llama-Nemo, absent for Jina (vision), and on average across the benchmark GQR consistently improves retrieval quality. We observe similar patterns for the $\mathrm{Recall}@5$ metric (Table 10).
On ViDoRe 1 (Table 5 in the Appendix) GQR is generally on par with the base models, neither harming performance nor providing a marked benefit. Notably, the results show that the benchmark suffers from saturation, with many subset scores reaching $90$ or higher (and indeed this was the direct motivation for the release of ViDoRe 2, Macé et al., 2025).
Comparison to hybrid baselines.

### v2

@@ Abstract
Multimodal encoders have pushed the boundaries of visual document retrieval, matching textual query tokens directly to image patches and achieving state-of-the-art performance on public benchmarks. Recent models relying on this paradigm have massively scaled the sizes of their query and document representations, presenting obstacles to deployment and scalability in real-world pipelines. Furthermore, purely vision-centric approaches may be constrained by the inherent modality gap still exhibited by modern vision-language models. In this work, we connect these challenges to the paradigm of hybrid retrieval, investigating whether a lightweight dense text retriever can enhance a stronger vision-centric model. Existing hybrid methods, which rely on coarse-grained fusion of ranks or scores, fail to exploit the rich interactions within each model’s representation space. To address this, we introduce Guided Query Refinement (GQR), a novel test-time optimization method that refines a primary retriever’s query embedding using guidance from a complementary retriever’s scores. Through extensive experiments on visual document retrieval benchmarks, we demonstrate that GQR allows vision-centric models to match the performance of models with significantly larger representations, while being up to 14x faster and requiring 54x less memory. Our findings show that GQR effectively pushes the Pareto frontier for performance and efficiency in multimodal retrieval. We release our code here.
@@ 1 Introduction
Visual document retrieval is the task of returning relevant documents – typically PDFs containing figures, tables, and other visual elements – in response to a textual query (mathew2021docvqadatasetvqadocument; mathew2021infographicvqa; li-etal-2024-multimodal-arxiv; zhu2022towards; faysse2025colpali). To tackle this task, neural retrieval pipelines often follow a text-centric approach, relying on OCR or vision-language models to convert source documents into textual chunks, and then constructing an index using semantic text encoders (karpukhin-etal-2020-dense; VisualMRC2021). An alternative, vision-centric, approach relies instead on multimodal encoder models. Building on the ColBERT (khattab2020colbert) late-interaction approach, ColPali-based (faysse2025colpali) encoders operate directly on image patches, and yield multi-vector embedding representations of images and queries.
@@
While this approach achieves state-of-the-art results on public benchmarks of visual document retrieval [^1] https://huggingface.co/blog/manu/vidore-v2 (mace2025vidore), open challenges within this paradigm remain. First, to pursue state-of-the-art performance, recent late-interaction multimodal retrievers [^2] Henceforth, we use “retrievers” and “encoders” interchangeably. For additional background on these emerging retrieval paradigms, refer to Appendix A. massively scale the length and dimensionality of query and document representations. This can incur a substantial latency and storage overhead, hindering the ability to provide an efficient and scalable solution. For example, Llama-NemoRetriever-Colembed-3B represents each document page with 10 MB of memory (xu2025llama_nemo), three orders of magnitude more than single-vector dense retrievers (Table 4). Secondly, a vision-centric approach for matching textual queries to textually rich documents may be limited by the substantial modality gap (clavié2025readbenchmeasuringdensetext; li2025closingmodalitygapmixed; role2025gapquantifyingreducingmodality) exhibited by modern vision-language models. These gaps motivate exploring complementary approaches for improving the performance of multimodal encoders.
@@
An early concept in the application of neural retrievers has been that of hybrid retrieval (dos-santos-etal-2015-learning; kuzi2020leveraging), where the outputs of different retrievers are aggregated at the level of ranks or query-document similarity scores (see Figure 1) to obtain the final list of retrieved documents. Hybrid retrieval (or hybrid search) most commonly refers to the combination of a neural semantic text retriever with a sparse lexical representation (e.g., BM25). More broadly, it reflects the notion that models relying on different types of representations can capture complementary aspects of the data, which can in turn be leveraged to boost the overall performance on the task.
In this work, we seek to connect these two threads, testing whether the paradigm of hybrid retrieval can complement modern multimodal encoders. Dense text retrievers typically have low online latencies and incur a small storage footprint, and they provide a uni-modal signal between the query and the document representation. Thus, hybrid retrieval between text and image encoders emerges as a natural candidate for enhancing the performance of a multimodal retrieval system. However, in this work we argue that standard hybrid retrieval methods rely on a rather coarse-grained view of the perspective of each retriever – they cannot utilize the rich query-document interactions within the model representation space (Figure 1).
@@
Aiming to harness this untapped potential of hybrid retrieval, in this work we propose Guided Query Refinement (GQR) (Figure 2), a novel approach for aggregating retriever outputs. Given a query at test time, GQR iteratively optimizes the query representations of a primary retriever with gradient descent, using similarity scores from a complementary retriever. The refined query representation softly incorporates the complementary retriever’s signal, remaining subject to the query–document interactions in the primary retriever space. This updated query embedding is then used to score the documents and return an updated document list. Notably, GQR is architecture-agnostic and can be used to optimize both single- and multi-vector query embeddings.
@@ 2 Methods
We begin this section by outlining prominent methods at each level (§2.1, §2.2). We explain that while early representation-level aggregation could be desired due to its richness, it is difficult to achieve in practice. We then present our method, Guided Query Refinement (GQR), which lies between the levels of scores and of representations (§2.3) and provide the motivation for it (§2.4).
@@ 2.1 Ranking-Level Aggregation
RRF (36196) combines ranked lists by weighting each item based on the reciprocal of its rank. The RRF constant $\kappa>0$ dampens the impact of very high ranks and controls how much credit is given to mid-list occurrences.
@@ 2.2 Score-Level Aggregation
Score-level aggregation operates one step deeper than ranking aggregation, operating on the real-valued similarity scores $s_{m}(q,d_{i})$ between the query and documents. To ensure that the scores of different retrievers are in the same scale and range, the common practice (bruch2023analysis) is to first apply a normalization function $N_{m}$ – yielding $\tilde{s}_{m}(q,d_{i})$ – and then aggregate across retrievers:
@@
More generally, the above methods can be viewed as a weighted aggregation, where the examples above are the uniform case, with each retriever assigned a weight $\alpha=1/M$ (for two retrievers, $\alpha=0.5$ for each). Given a development set, these weights can be fit (bruch2023analysis). Here, for $M=2$, the two retrievers are assigned relative weights $\alpha$ and $1-\alpha$, yielding the parameterized variants Average Ranking - Tuned, RRF - Tuned, Score Aggregation (Min-Max) - Tuned, and Score Aggregation (SoftMax) - Tuned. Details are in Appendix C.
@@ 2.3 Guided Query Refinement
Representation-level information carries the richest potential for effective aggregation. Embedding-level projections that align representational spaces are used extensively in modern vision–language systems to combine visual and textual inputs (radford2021clip; jia2021align; li2021albef; li2022blip). At test time, however, operating directly on representations is hindered by heterogeneity: encoders may use a single vector or many vectors per document and query, and they operate within differing dimensionalities and scales. Thus, with strict latency and memory budgets and without access to supervision, aggregation at this level is not trivial. [^4] Concatenating single-vector embeddings is feasible, yet under dot-product scoring this reduces to an unnormalized sum of separate scores, and does not enable interaction between the spaces.
Our goal in this work is to exploit the rich information in query and document representations while remaining architecture-agnostic, lightweight, and practical. To this end, we propose Guided Query Refinement (GQR), a novel method for combining the outputs of two retrievers – a primary retriever $m_{1}$ and a complementary retriever $m_{2}$ (Algorithm 1, Figure 2). GQR refines $m_{1}$’s query representation based on the signal of $m_{2}$’s scores.
@@
We denote the initial query embedding of $m_{1}$ by $z^{(0)}=e_{1}^{q}$ (GQR is applicable to both single and multi-vector embeddings), and we update it at each step $t$, $z^{(t)}$ for $T$ steps. At step $t$, the consensus distribution of $m_{1}$ and $m_{2}$ is defined as
@@
Here, $\mathrm{KL}$ is the Kullback–Leibler divergence (Equation 12).
We apply a gradient step on the query representation with step size $\alpha$ [^5] We define GQR with gradient descent for simplicity, but in practice we found Adam (kingma2015adam) to perform better and use it as the optimizer.,
@@
We then compute the final scores from retriever $m_{1}$,
@@
Figure 2: Guided Query Refinement (GQR). Stage 1: Two retrievers independently encode the query and retrieve top-$K$ documents, forming a candidate pool. Stage 2: The primary query embedding is iteratively refined ($z^{(t)}$) over $T$ iterations, by minimizing the KL divergence between a consensus distribution and the primary distribution.
2.4 GQR - Motivation and Rationale
Our approach is inspired by test time optimization methods that rely on pseudo-relevance feedback from a stronger cross-encoder (yu2021improvingqueryrepresentationsdense; sung-etal-2023-optimizing; 10.1145/3726302.3730160). Minimizing $\mathrm{KL}$ pushes $p_{1}(\cdot\mid z^{(t)})$ to place higher probability where the other distribution does, and to reduce probability where the other distribution is low. However, instead of relying on a heavy cross-encoder, here we utilize feedback from a lightweight bi-encoder whose performance can be on par or even weaker than the primary encoder, and thus the consensus distribution is set as $p_{\mathrm{avg}}$.
Compared to a simple weighted average of scores, GQR operates by updating the query representation of $m_{1}$ rather than defining a new scoring rule. The complementary distribution $p_{\mathrm{avg}}^{(t)}$ is used as a learning signal, and the gradient of $\mathcal{L}^{(t)}$ is computed through $s_{1}(z,d)$, so any change in the ranking must be achievable by moving the query within the primary model’s embedding space. Intuitively, this allows feedback from $m_{2}$ to influence the query in a way that is constrained by $m_{1}$’s notion of similarity, since documents contribute to the update through their embeddings and their relative probabilities. In score-level fusion, each document’s probability is pulled toward the secondary retriever’s distribution by the same fixed amount, so the update is always a uniform weighted average. In GQR, different documents can shift by different magnitudes along a non-linear trajectory dictated by the geometry of that space. In cases where $m_{2}$ is weaker or misaligned, this setup can help incorporate its signal more softly, since it is filtered through $m_{1}$’s representation rather than directly overriding scores.
@@ 3.1 Setup
Visual document retrieval (mathew2021docvqadatasetvqadocument; mathew2021infographicvqa; li-etal-2024-multimodal-arxiv; zhu2022towards; faysse2025colpali) assumes a corpus of documents, that contain visual elements such as charts, images, and tables, and a set of document-grounded textual queries. The goal is to retrieve the most relevant documents for each query. We conduct experiments on ViDoRe 1 (faysse2025colpali), ViDoRe 2 (mace2025vidore) and the ViDoRe 3 benchmarks, which are established benchmarks for this task. Corpus documents are embedded by encoder models, either directly from page images, or following ingestion of document pages into text (see Appendix D).
@@
We evaluate a diverse pool of multimodal and textual state-of-the-art retrieval models. The Colpali-based set includes three encoders: Colnomic-Embed-Multimodal-7B (nomicembedmultimodal2025), Jina-Embeddings-v4 (gunther2025jina_v4), and Llama-NemoRetriever-Colembed-3B (xu2025llama_nemo). The set of text models includes Linq-Embed-Mistral (choi2024linq) and Qwen3-Embedding-4B (qwen3embedding), as well as Jina-Embeddings-v4 in its multi-vector textual configuration. This yields $3$ text-based models and $3$ image-based models in total. See Table 4 for details on the models. For the ViDoRe 3 benchmark, we excluded Llama-NemoRetriever-Colembed-3B due to its high computational overhead, which makes evaluation impractical over this larger benchmark.
@@
We use $\mathrm{NDCG}@5$ as the primary metric for our evaluations (we also report $\mathrm{Recall}@5$ in Appendix F.) For each ColPali-based vision-centric model, we test each of the text-centric models as the complementary retriever used for GQR. This yields $9$ GQR pairs in total for ViDoRe 1 and ViDoRe 2 and $6$ for ViDoRe 3, $3$ per ColPali-based model. We also evaluate $4$ different hybrid methods for each vision-text model pair. See Appendix D for technical details about tuning the hyperparameters for ViDoRe 1 and ViDoRe 2.
To simulate real production use cases, we evaluate both tuned and zero-shot settings. We begin with the tuned setting on ViDoRe 1 and ViDoRe 2, and then use the findings from ViDoRe 2 to guide a zero-shot evaluation on ViDoRe 3.
3.2 Comparison to Single-Retriever Pipelines
@@
Results with Hyperparameter Tuning.
@@
Table 1 reports GQR against the corresponding models on ViDoRe 2. The first block lists the text-only models, which average between $46.8$ for Qwen and $55.3$ for Linq. In each subsequent block, a ColPali-based retriever is fixed, and we show its score alongside the GQR variants, with deltas computed relative to the primary retriever. For Colnomic-7B, the average score rises from $60.3$ to $63.1$ with query refinement from Jina (text) $(+2.8)$ and to $62.8$ with Linq-Embed $(+2.5)$. For Llama-Nemo, the strongest model on ViDoRe 2 to date, GQR improves the average from $63.0$ to $65.2$ with Linq-Embed $(+2.2)$. Notably, the text models clearly underperform the ColPali-based retrievers, yet with GQR the complementary signal they provide boosts performance. This is clearest in the Llama-Nemo versus Qwen3 setting, where GQR delivers a small gain of $\uparrow$+0.3+0.3, despite a $16.2$ point gap in base $\mathrm{NDCG}@5$ (Llama-Nemo $63.0$ vs. Qwen $46.8$). Subsets where GQR harms performance are rare for Colnomic-7B and Llama-Nemo, absent for Jina (vision), and on average across the benchmark GQR consistently improves retrieval quality. We observe similar patterns for the $\mathrm{Recall}@5$ metric (Table 12). On ViDoRe 1 (Table 7 in the Appendix) GQR is generally on par with the base models, yet the benchmark clearly suffers from saturation, with many subset scores reaching $90$ or higher.
Table 2: $\mathrm{NDCG}@5$ over ViDoRe 3, by primary and complementary models. Columns show subset scores, average, and deltas relative to the no-refinement baseline within each primary model.
Primary Model GQR complementary model Avg cs energy fin_en fin_fr hr industrial pharma physics
val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$
Jina (text) $51.6$ $\uparrow$+0.0 $66.6$ $\uparrow$+0.0 $60.2$ $\uparrow$+0.0 $54.0$ $\uparrow$+0.0 $40.1$ $\uparrow$+0.0 $54.5$ $\uparrow$+0.0 $42.2$ $\uparrow$+0.0 $55.3$ $\uparrow$+0.0 $40.5$ $\uparrow$+0.0
Linq-Embed $44.2$ $\uparrow$+0.0 $64.6$ $\uparrow$+0.0 $48.5$ $\uparrow$+0.0 $40.1$ $\uparrow$+0.0 $28.7$ $\uparrow$+0.0 $42.6$ $\uparrow$+0.0 $30.1$ $\uparrow$+0.0 $57.7$ $\uparrow$+0.0 $41.4$ $\uparrow$+0.0
Qwen3 $46.0$ $\uparrow$+0.0 $67.6$ $\uparrow$+0.0 $50.1$ $\uparrow$+0.0 $38.9$ $\uparrow$+0.0 $27.3$ $\uparrow$+0.0 $44.6$ $\uparrow$+0.0 $37.8$ $\uparrow$+0.0 $58.1$ $\uparrow$+0.0 $43.9$ $\uparrow$+0.0
Colnomic-7B
No refinement $55.7$ $\uparrow$+0.0 $74.0$ $\uparrow$+0.0 $62.3$ $\uparrow$+0.0 $54.2$ $\uparrow$+0.0 $42.7$ $\uparrow$+0.0 $57.1$ $\uparrow$+0.0 $49.5$ $\uparrow$+0.0 $62.0$ $\uparrow$+0.0 $44.1$ $\uparrow$+0.0
Jina (text) 57.4 $\uparrow$+1.7 $74.8$ $\uparrow$+0.8 64.1 $\uparrow$+1.8 58.6 $\uparrow$+4.4 44.6 $\uparrow$+1.9 59.4 $\uparrow$+2.3 50.8 $\uparrow$+1.3 $62.8$ $\uparrow$+0.8 $44.5$ $\uparrow$+0.4
Linq-Embed $57.1$ $\uparrow$+1.4 $75.1$ $\uparrow$+1.1 $63.4$ $\uparrow$+1.1 $56.9$ $\uparrow$+2.7 $43.9$ $\uparrow$+1.2 $58.1$ $\uparrow$+1.0 50.8 $\uparrow$+1.3 63.4 $\uparrow$+1.4 45.7 $\uparrow$+1.6
Qwen3 $56.7$ $\uparrow$+1.0 75.2 $\uparrow$+1.2 $63.1$ $\uparrow$+0.8 $56.0$ $\uparrow$+1.8 $42.6$ $\downarrow$-0.1 $58.0$ $\uparrow$+0.9 $50.5$ $\uparrow$+1.0 $63.1$ $\uparrow$+1.1 $45.4$ $\uparrow$+1.3
Jina (vision)
No refinement $53.4$ $\uparrow$+0.0 $69.2$ $\uparrow$+0.0 $59.4$ $\uparrow$+0.0 $52.4$ $\uparrow$+0.0 $39.8$ $\uparrow$+0.0 $54.7$ $\uparrow$+0.0 $47.8$ $\uparrow$+0.0 $61.2$ $\uparrow$+0.0 $43.0$ $\uparrow$+0.0
Jina (text) $55.1$ $\uparrow$+1.7 $70.2$ $\uparrow$+1.0 61.5 $\uparrow$+2.1 56.3 $\uparrow$+3.9 42.6 $\uparrow$+2.8 57.1 $\uparrow$+2.4 $48.1$ $\uparrow$+0.3 $61.6$ $\uparrow$+0.4 $43.5$ $\uparrow$+0.5
Linq-Embed 55.4 $\uparrow$+2.0 $72.0$ $\uparrow$+2.8 $61.2$ $\uparrow$+1.8 $54.9$ $\uparrow$+2.5 $41.5$ $\uparrow$+1.7 $56.6$ $\uparrow$+1.9 $48.6$ $\uparrow$+0.8 63.8 $\uparrow$+2.6 $45.1$ $\uparrow$+2.1
Qwen3 $55.0$ $\uparrow$+1.6 72.3 $\uparrow$+3.1 $60.4$ $\uparrow$+1.0 $54.1$ $\uparrow$+1.7 $39.7$ $\downarrow$-0.1 $55.9$ $\uparrow$+1.2 49.3 $\uparrow$+1.5 $63.1$ $\uparrow$+1.9 45.6 $\uparrow$+2.6
Results - Zero Shot.
For ViDoRe 3, we use a single hyperparameter configuration across all subsets and model pairs, a choice informed by our hyperparameter investigation on top of ViDoRe 2 in subsection 4.1. Table 2 shows that GQR improves performance across all model pairs and benchmark subsets, consistently outperforming the no-refinement baseline. For example, on Colnomic-7B it raises the average $\mathrm{NDCG}@5$ from $55.7$ to $57.4$ with Jina as the complementary model (a gain of +1.7+1.7), and delivers per-subset gains as high as +4.4+4.4 on fin_en. Similar improvements appear for Jina (vision), where GQR increases the average from $53.4$ to $55.4$ with Linq-Embed (+2.0+2.0).
3.3 Comparison to hybrid retrieval pipelines